I am supposed to write a quadratic equation in the variable x having the given numbers as a solution. I am supposed to type the equation in standard form, ax^2 + bx + c = 0. The solutions are -5,1
Is the answer x^2 - 25 = 0? Answer by stanbon(75887) 
 I am supposed to write a quadratic equation in the variable x having the given numbers as a solution. I am supposed to type the equation in standard form, ax^2 + bx + c = 0. The solutions are -5,1
-------------------
Since -5 is a solution, x+5 is a factor.
----
Since 1 is a solution, x-1 is a factor.
----
Quadratic: f(x) = (x+5)(x-1) = x^2+4x-5
===========================================
Cheers,
Stan H.